How does the length of a rectangle change when the width is decreased by 6 units and its perimeter remains the same ?
Answer by ad_alta(240)

Let the perimeter be 'P' and the length and width 'l' and 'w' respectively. We know 2w+2l=P. We also know that l' (the new length) satisfies 2l'+2(w-6)=P=2l+2w. This means that l'=l+6. This is sort of obvious, but the new length must be 6 units longer [but now you've actually proved that result!].